OpenStudy (anonymous):

Name the vertex, domain and range, and the y-intercept of f(x) = x^2 + 6x + 4

OpenStudy (anonymous):

y-intercept is when x = 0, so substitute 0 for x. domain is the same as it is for all quadratics. range depends on vertex... So the vertex: complete the square: y = x^2 + 6x + 4 ...

OpenStudy (anonymous):

x^2 + 6x + 9 is equal to (x + 3)^2, so add 5 and subtract 5: y = x^2 + 6x + 4 +5 -5 y = x^2 + 6x + 9 - 5 y = (x+3)^2 -5 vertex form! Can you get it from here?

OpenStudy (anonymous):

i got (-3,-8) is that right?

OpenStudy (anonymous):

-3,-5